OpenStudy (anonymous):

Hi can you help me with 2c and 2d not the rest. If you can show me the proper steps for one i shall attempt the second. Thannnks

OpenStudy (anonymous):

I'll help through the steps for 2c if someone would like 2d.

OpenStudy (anonymous):

i will do 2d

OpenStudy (anonymous):

Alright pretty sure I did it right, its been a while since calc 1 xD dy/dx = 2(sqrt(x) - cotx) + 2x((1/2)x^(-.5) - (cscx)^2) = \[3\sqrt{x} - 2\cot x - 2x \csc ^{2}x\]

OpenStudy (anonymous):

\[2\sqrt{x} - cotx + 2x (1/2)x ^{-1/2} - (\csc x)^{2 }\] is this what u meant? thanks